OpenStudy (hackergod):

Math help anyone ? 9th grade!!! Algebra 1

OpenStudy (anonymous):

I'll see if i can help :)

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

Wht school do u go too

OpenStudy (hackergod):

Thanks :) 2. Which equation has (2,-6) and (-3,4) as solutions? A. y=1/2x-7 B.y=2x-10 C.y=-1/2x-5 or D. y=-2x-2

OpenStudy (anonymous):

And hold on :P

OpenStudy (anonymous):

Do you want me to help you on every question ?

OpenStudy (anonymous):

Its D.y=-2x-2

OpenStudy (hackergod):

Do you know 6 or 7?

OpenStudy (anonymous):

Number seven is "C"

OpenStudy (anonymous):

And i'm pretty sure number six is "A"

OpenStudy (hackergod):

Thank you so much!!

OpenStudy (anonymous):

Tell me if you get them right :)

OpenStudy (hackergod):

I will once my teacher grades them :)

OpenStudy (anonymous):

Okay :P